CO/He GLOW DISCHARGE TEMPERATURE MEASUREMENTS USING CO 4th POSITIVE BAND LIF

\underline{R. J. LEIWEKE} and W. R. LEMPERT, Nonequilibrium Thermodynamics Laboratory, Department of Mechanical Engineering and Department of Chemistry, The Ohio satte University, Columbus, OH 43210.